OpenStudy (anonymous):

Write the expression using scientific notation - (2.5 * 10^4) (4 * 10^3) A. 10 * 10^7 B. 10 * 10^8 C. 1 * 10^8 D. 10 * 10^12

OpenStudy (ajprincess):

what is 2.5*4=?

OpenStudy (anonymous):

10

OpenStudy (ajprincess):

yup good:) \[a^b*a^c=a^{b+c}\] nw \[10^4*10^3=?\]

OpenStudy (anonymous):

Hmm.. I wasn't sure if I added the exponents, or multiplied them... so in that case it'd be 10 * 10^7

OpenStudy (ajprincess):

only adding them gives u 7. so \[10^4*10^3=10^7\] \[10=1*10^1\] \[10^1*10^7=?\]

OpenStudy (ajprincess):

usually numbers written in scientific notation will be in the form 2*10^5 2.6^10^7